# Jarník-Besicovitch type theorems for semisimple algebraic groups

### v1

@@ 1.2. Main results
Let $\rho:\operatorname{\mathbf{SL}}_{n}\to\operatorname{\mathbf{GL}}(V)$ be the standard representation of $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$ on the complex vector space $V=\mathbb{C}^{n}$ defined by
$\rho(g)\cdot v=g\cdot v\quad(g\in\operatorname{\mathbf{SL}}_{n},v\in V)$
via matrix multiplication. Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$, $\beta_{0}$ the highest weight of $\rho$ with respect to $\{a_{t}\}_{t\in\mathbb{R}}$ defined as in Theorem 1.3, and $\nu_{0}$ the $\mathbb{Q}$-root defined as in Theorem 1.4. Then for any $0\leq\tau<\beta_{0}(a_{-1})$, we have
$\dim_{H}S_{\tau}^{c}=\dim\mathbf{G}-\frac{n\cdot\tau}{\nu_{0}(a_{1})}.$
The second result generalizes the main results in [17].
@@
Let $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ where $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$, $V=\mathfrak{g}=\mathfrak{sl}_{n}$ is the Lie algebra of $\operatorname{\mathbf{SL}}_{n}$ and $\rho=\operatorname{Ad}$ is the adjoint representation of $\mathbf{G}$. Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter diagonal subgroup in $\mathbf{G}$, and $\nu_{0}$ the $\mathbb{Q}$-root defined as in Theorem 1.4. Then for any $0\leq\tau<\nu_{0}(a_{1})$, we have
$\dim_{H}S_{\tau}^{c}=\dim\mathbf{G}-\frac{(n-1)\cdot\tau}{\nu_{0}(a_{1})}.$
We also deduce the following
Theorem 1.8.
Let $\rho:\operatorname{\mathbf{SL}}_{2}\to\operatorname{\mathbf{GL}}(V)$ be an irreducible representation of $\operatorname{\mathbf{SL}}_{2}$ defined over $\mathbb{Q}$. Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{2}(\mathbb{R})$, $\beta_{0}$ the highest weight of $\rho$ with respect to $\{a_{t}\}_{t\in\mathbb{R}}$ defined as in Theorem 1.3, and $\nu_{0}$ the $\mathbb{Q}$-root in $\operatorname{\mathbf{SL}}_{2}$ defined as in Theorem 1.4. Then for any $0\leq\tau<\beta_{0}(a_{-1})$, we have
$\dim_{H}S_{\tau}^{c}=3-\frac{\tau}{\beta_{0}(a_{-1})}.$
Remark 1.9.
Let $\rho:\operatorname{\mathbf{SL}}_{2}\to\operatorname{\mathbf{GL}}(V)$ be the irreducible representation in Theorem 1.8 with $\dim V=n+1$ $(n\geq 1)$, $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{2}(\mathbb{R})$ and let $\rho_{0}$ be the standard representation of $\operatorname{\mathbf{SL}}_{n+1}$ on $\mathbb{C}^{n+1}$. Note that $V\cong\mathbb{C}^{n+1}$ and $\rho(\operatorname{\mathbf{SL}}_{2})\subset\operatorname{\mathbf{SL}}_{n+1}\subset\operatorname{\mathbf{GL}}(V)$. Then by Theorem 1.8, we have
$\dim_{H}S_{\tau}(\rho_{0},\{\rho(a_{t})\}_{t\in\mathbb{R}})^{c}\cap\rho(\operatorname{\mathbf{SL}}_{2}(\mathbb{R}))=\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}=3-\frac{\tau}{\beta_{0}(a_{-1})}$
where $\beta_{0}$ is the highest weight of $\rho$ with respect to $\{a_{t}\}_{t\in\mathbb{R}}$ defined as in Theorem 1.3. On the other hand, by Theorem 1.6
$\dim_{H}S_{\tau}(\rho_{0},\{\rho(a_{t})\}_{t\in\mathbb{R}})^{c}=\dim\operatorname{\mathbf{SL}}_{n+1}-\frac{(n+1)\cdot\tau}{\nu_{0}(\rho(a_{1}))}$
where $\nu_{0}$ is the $\mathbb{Q}$-root in $\operatorname{\mathbf{SL}}_{n+1}$ with respect to $\{\rho(a_{t})\}_{t\in\mathbb{R}}$ defined as in Theorem 1.4. This gives an example of Jarník-Besicovitch type theorems for some algebraic subvarieties in affine spaces.
Remark 1.10.
One can see from the arguments in this note that all the theorems stated above also work for any open bounded subset $U$ in $\mathbf{G}(\mathbb{R})$, i.e. $\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}\cap U=\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}$. (See Remark 5.6.)
1.3. Discussions and open problems
[…]
$\liminf_{t\to\infty}-\frac{1}{t}\log\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})=\tau.$
When $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$ and $\rho$ is the standard representation of $\mathbf{G}$, the Hausdorff dimension and the packing dimension of $\operatorname{Sing}(\tau)$ are discussed in [43] in the case that $\{a_{t}\}_{t\in\mathbb{R}}$ is a singular one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$, and the proofs rely on a variational principle in the parametric geometry of numbers. This variational principle is later generalized in [39] for any one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{n}(\mathbb{R})$.
For the lower bound and the upper bound we obtain in Theorem 1.3 and Theorem 1.4, we have computed in several cases (e.g. Theorems 1.61.7 and 1.8) that these two bounds are equal. It would be interesting to compute these two bounds in other cases and determine the representations $\rho$ for which these two bounds are equal.
@@ 2. Preliminaries
In this section, we list some preliminaries needed in this note. Then we give a definition of rational elements in $\mathbf{G}(\mathbb{R})$ and discuss some properties of rational elements.
We first need the reduction theory of arithmetic subgroups of $\mathbf{G}(\mathbb{R})$ [5]. Let $K$ be a maximal compact subgroup in $\mathbf{G}(\mathbb{R})$ and $\Gamma$ an arithmetic subgroup in $\mathbf{G}(\mathbb{Z})$. We fix a minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ in $\mathbf{G}$ containing $\mathbf{T}$. Denote by $\Phi$ the set of $\mathbb{Q}$-roots with respect to $\mathbf{T}$, $\Phi^{+}$ the set of positive $\mathbb{Q}$-roots corresponding to the minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ and $\Delta$ the set of simple $\mathbb{Q}$-roots in $\Phi^{+}$. For any parabolic $\mathbb{Q}$-subgroup $\mathbf{P}$, denote by $R_{u}(\mathbf{P})$ its unipotent radical. Let $M$ be the connected component of identity in the unique maximal $\mathbb{Q}$-anisotropic subgroup in $Z_{\mathbf{G}(\mathbb{R})}(\mathbf{T}(\mathbb{R}))$ (= the centralizer of $\mathbf{T}(\mathbb{R})$ in $\mathbf{G}(\mathbb{R})$). For $\eta>0$, denote by
@@
A Siegel set in $\mathbf{G}(\mathbb{R})$ is a subset of the form $S_{\eta,\Omega}=K\cdot T_{\eta}\cdot\Omega$ for some $\eta>0$ and a relatively compact open subset $\Omega$ containing identity in $M\cdot R_{u}(\mathbf{P}_{0})(\mathbb{R})$, and the group $\mathbf{G}(\mathbb{R})$ can be written as
@@
for some Siegel set $S_{\eta,\Omega}$ and some finite subset $\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$. Moreover, the finite set $\mathcal{K}$ satisfies the property that
@@
where $\mathbf{P}_{0}$ is the minimal parabolic $\mathbb{Q}$-subgroup. Denote by
@@
and we may assume that $e\in\mathcal{K}$. In what follows, we choose $\Gamma$ to be an arithmetic subgroup in $\mathbf{G}(\mathbb{Z})\cap\mathbf{G}(\mathbb{R})^{0}$, where $\mathbf{G}(\mathbb{R})^{0}$ denotes the connected component of identity in the Lie group $\mathbf{G}(\mathbb{R})$, so that $\rho(\Gamma)$ preserves the lattice $\mathbb{Z}^{d}$ in $V$. Without loss of generality, we may assume that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\mathbf{P}_{0})$ and its unstable horospherical subgroup is contained in $R_{u}(\overline{\mathbf{P}}_{0})$ where $\overline{\mathbf{P}}_{0}$ is the opposite minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{0}$ determined by $\Phi\setminus\Phi^{+}$. We also write $\mathbf{P}_{0}=\mathbf{M}_{0}\cdot R_{u}(\mathbf{P}_{0})$ where $\mathbf{M}_{0}$ is the centralizer of $\mathbf{T}$ in $\mathbf{G}$, and write $\mathbf{M}_{a}$ for the maximal $\mathbb{Q}$-anisotropic subgroup in $\mathbf{M}_{0}$ so that $\mathbf{M}_{0}=\mathbf{T}\cdot\mathbf{M}_{a}$.
Now we choose a maximal $\mathbb{Q}$-torus $\mathbf{S}$ in $\mathbf{P}_{0}$ containing $\mathbf{T}$. The Lie algebra $\mathfrak{g}$ of $\mathbf{G}$ can be written as a direct sum
$\mathfrak{g}=\mathfrak{g}_{0}\oplus\bigoplus_{\alpha\in\Psi}\mathfrak{g}_{\alpha}$
of root spaces relative to $\mathbf{S}$ via the adjoint representation of $\mathbf{G}$, where $\Psi$ denotes the set of roots in $\mathfrak{g}$ relative to $\mathbf{S}$. We may determine a set of positive roots in $\Psi$, which we denote by $\Psi^{+}$, such that
@@
The set of simple roots in $\Psi^{+}$ is denoted by $\Pi$.
It is known that there are complete classifications of irreducible representations of complex semisimple groups and Lie algebras [21, 27]. Let $V=\bigoplus_{\lambda}V_{\lambda}$ be the decomposition of the vector space $V$ in $\rho$ into the direct sum of weight spaces $V_{\lambda}$ relative to $\mathbf{S}$. According to the theorem of highest weight, there is a unique highest weight $\lambda_{0}$ among the weights $\lambda$’s (the order is determined by $(\Psi,\Psi^{+},\Pi)$) such that

### v2

@@ 1.3.1. Connection to Diophantine approximation on flag varieties
Let $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a decreasing function. Then $0\leq\gamma(\psi)\leq\infty$. If the highest weight $\chi$ of $\rho_{\chi}$ is a multiple of $\sum_{\alpha\in\Phi(R_{u}(\mathbf{P}))}{\alpha}$, then the Hausdorff dimension of $E_{\mathbf{X}}(\psi)^{c}$ (with respect to a standard Riemannian metric on $\mathbf{X}(\mathbb{R})$) equals
$\dim_{H}E_{\mathbf{X}}(\psi)^{c}=$ $\dim\mathbf{X}-\left(\frac{1}{\beta_{\chi}}-\frac{1}{\gamma(\psi)+\beta_{\chi}}\right)\frac{1}{\chi(a_{-1})\nu_{0}(a_{-1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\mathbf{P}))}\alpha(a_{-1})$
for any $0\leq\gamma(\psi)\leq\infty$.
1.3.2. Connection to rational approximation to linear subspaces
Diophantine approximation on the Grassmann variety of $l$-dimensional subspaces in $\mathbb{R}^{n}$ by $k$-dimensional $\mathbb{Q}$-subspaces $(1\leq k\leq l<n)$ is a problem suggested by Schmidt in [71]. It has attracted increasing attention and one may refer to e.g. [67, 19, 13, 43, 44, 26]. Recently, de Saxcé has made a breakthrough in [16] which answers several Schmidt’s problems in [71]. Here we discuss a connection between Theorem 1.6 and [16]. We mainly follow the setup in [16].
For any $1\leq l\leq n-1$ ($l\in\mathbb{N}$), let $X_{l}(\mathbb{R})$ be the Grassmann variety of $l$-dimensional linear $\mathbb{R}$-subspaces in $\mathbb{R}^{n}$. Let $X_{l}(\mathbb{Q})$ be the set of $\mathbb{Q}$-rational points in $X_{l}(\mathbb{R})$, i.e. the set of $l$-dimensional $\mathbb{Q}$-subspaces in $\mathbb{R}^{n}$. We define a distance $d(\cdot,\cdot)$ between subspaces in $\mathbb{R}^{n}$ as follows. Let $P_{1}$ and $P_{2}$ be two linear $\mathbb{R}$-subspaces. If $P_{1}=\mathbb{R}\cdot v_{1}$ and $P_{2}=\mathbb{R}\cdot v_{2}$ are two lines $(v_{1},v_{2}\in\mathbb{R}^{n}\setminus\{0\})$, then the distance between $P_{1}$ and $P_{2}$ is the usual distance on the projective space $\mathbf{P}\mathbb{R}^{n-1}$, i.e.
$d(P_{1},P_{2})=\sin(\theta(P_{1},P_{2}))=\frac{\|v_{1}\wedge v_{2}\|}{\|v_{1}\|\|v_{2}\|}$
where $\theta(P_{1},P_{2})$ is the angle between $P_{1}$ and $P_{2}$, and the norms are the Euclidean norms on $\mathbb{R}^{n}$ and $\bigwedge^{2}\mathbb{R}^{n}$ respectively. If $P_{1}$ is a line, then the distance between $P_{1}$ and $P_{2}$ is defined by
$d(P_{1},P_{2})=\min\{d(P_{1},Q):Q\textup{ is a one-dimensional linear subspace in }P_{2}\}.$
If $P_{1}$ and $P_{2}$ are two general subspaces in $\mathbb{R}^{n}$, then we define the distance between $P_{1}$ and $P_{2}$ by
[…]
For $1\leq k\leq n-1$, we define a height function $H(\cdot)$ on $X_{k}(\mathbb{Q})$ as follows. For any $\mathbb{Q}$-subspace $P\in X_{k}(\mathbb{Q})$, we choose an integral basis $\{v_{i}\}_{1\leq i\leq k}\subset\mathbb{Z}^{n}$ for $P$ and define the height of $P$ by
$H(P)=\|v_{1}\wedge\cdots\wedge v_{k}\|$
where the norm is the Euclidean norm on $\bigwedge^{k}\mathbb{R}^{n}$.
Now let $1\leq k\leq l<n$. With the distance $d(\cdot,\cdot)$ and the height function $H(\cdot)$ as above, for any $x\in X_{l}(\mathbb{R})$, we may define the Diophantine exponent of $x$ for approximation by $k$-dimensional rational subspaces by
$\beta_{k}(x)=\inf\{\beta>0:d(v,x)\geq C\cdot H(v)^{-\beta}\;(\forall v\in X_{k}(\mathbb{Q}))\textup{ for some }C>0\}.$
Theorem 1.13 ([16, Theorem 1]).
Let $1\leq k\leq l<n$. For any $x\in X_{l}(\mathbb{R})$, we have
$\beta_{k}(x)\geq\frac{n}{k(n-l)}$
and the equality holds for almost every point $x\in X_{l}(\mathbb{R})$ with respect to a Riemannian volume form on $X_{l}(\mathbb{R})$.
Definition 1.14.
For any decreasing function $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$, define the set of $\psi$-Diophantine subspaces in $X_{l}(\mathbb{R})$ for approximation by $k$-dimensional $\mathbb{Q}$-subspaces by
$E_{l,k}(\psi)=\{x\in X_{l}(\mathbb{R}):d(v,x)\geq C\cdot H(v)^{-\frac{n}{k(n-l)}}\cdot\psi(H(v))\;(\forall v\in X_{k}(\mathbb{Q}))\textup{ for some }C>0\}.$
We denote by $E_{l,k}(\psi)^{c}$ the complement of $E_{l,k}(\psi)$ in $X_{l}(\mathbb{R})$.
As in §1.3.1, for any function $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$, define the lower order at infinity of $\psi$ by
$\gamma(\psi)=\liminf_{t\to\infty}\left(-\frac{\ln\psi(t)}{\ln t}\right).$
From a correspondence theorem [16, Proposition 1] and Theorem 1.6 with the representation $\rho_{k}:\operatorname{\mathbf{SL}}_{n}\to\operatorname{\mathbf{GL}}(V)$ $(V=\bigwedge^{k}\mathbb{C}^{n})$ which is the $k$-th exterior product of the standard representation of $\operatorname{\mathbf{SL}}_{n}$, we may deduce the following result.
@@ 1.3.2. Connection to rational approximation to linear subspaces
Let $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a decreasing function, and $1\leq k\leq l<n$. Then $0\leq\gamma(\psi)\leq\infty$, and the Hausdorff dimension of $E_{l,k}(\psi)^{c}$ (with respect to a standard Riemannian metric on $X_{l}(\mathbb{R})$) equals
$\dim_{H}E_{l,k}(\psi)^{c}=(l-k)(n-l)+\frac{n}{n/(k(n-l))+\gamma(\psi)}$
for any $0\leq\gamma(\psi)\leq\infty$.
1.4. Strategy of the proofs and organization of the paper
The investigation of the shrinking target problem in $\operatorname{\mathbf{SL}}_{3}(\mathbb{R})/\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})$ in [25] provides a strategy to address the main problem above. However, in this paper, since we are dealing with a shrinking target problem in a homogeneous space of a semisimple algebraic group $\mathbf{G}$ from the perspective of representations, the studies on the structure of the group $\mathbf{G}$ and the structure of the representation $\rho$ would be fundamental in our analysis. Due to the complexities of the structures of $\mathbf{G}$ and $\rho$, we have to upgrade the method in [25] considerably to tackle the main problem. Here we may reduce the problem to computing $\dim_{H}S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
We first define a notion of rational elements in $\mathbf{G}(\mathbb{R})$ according to the representation $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$, and mainly focus on the rational elements in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. The set of rational elements in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ is usually not countable, but a collection of leaves in a foliation in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. To get a lower bound of the Hausdorff dimension of $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$, we would like to choose neighborhoods around these (leaves of) rational elements and construct a Cantor-type subset. In order to apply Hausdorff dimension formulas, at each level of the Cantor-type subset, the neighborhoods should be disjoint. In our case, it may happen that the neighborhoods at each level are not disjoint if we proceed in the usual way. It implies that there are surplus rational elements in the selected leafs, and we have to sieve them out so that the neighborhoods around the remaining rational elements are disjoint. It is also required that the proportion of the remaining rational elements is not small compared to the set of all rational elements in order to avoid any loss of Hausdorff dimension. For this purpose, we will need the mixing property of the flow $\{a_{t}\}_{t\in\mathbb{R}}$. The mixing property of $\{a_{t}\}_{t\in\mathbb{R}}$, together with the structure of the Siegel sets in $\mathbf{G}(\mathbb{R})$, not only helps us sieve out surplus rational elements, but also gives an asymptotic estimate about the measure of the subset of the remaining rational elements. Then we may apply the Hausdorff dimension formula of a Cantor-type set and get a lower bound for $\dim_{H}S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
To get an upper bound of $\dim_{H}S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$, we will need the Bruhat decomposition of $\mathbf{G}$, or precisely the decomposition
$\mathbf{G}(\mathbb{Q})=\bigcup_{w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q}).$
For any $p\in S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$, from the structures of the representation $\rho$ and the Siegel sets in $\mathbf{G}(\mathbb{R})$, one may deduce that there exist a Weyl element $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, a sequence of rational elements $\{q_{k}\}_{k\in\mathbb{N}}$ and a sequence of subsets $\{E_{k}\}_{k\in\mathbb{N}}$ in $\mathbf{F}_{w}(\mathbb{R})$ (defined from the preimages of certain open subsets under the morphism $\Psi_{w}$ contracted by the action of $\{a_{t}\}_{t\in\mathbb{R}}$) such that
$p\in E_{k}\cdot w^{-1}q_{k}\;(k\in\mathbb{N}).$
So we can divide the set $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ into finitely many subsets $\mathcal{E}_{w,\epsilon}(\psi)$ (indexed by the elements $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$), each of which corresponds to a type of Diophantine approximation in the sense that any $p\in\mathcal{E}_{w,\epsilon}(\psi)$ is covered by infinitely many subsets of the form $E\cdot w^{-1}q$ described as above. Now fix $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$. As mentioned before, the set of rational elements $q$ is a collection of leaves in some foliation in $\mathbf{G}(\mathbb{R})$, and by the structure of the representation $\rho$ and the morphism $\Psi_{w}$, the shape of the subset $E$ defined above may be arbitrary and there may be rational elements $q$ far away from a point $p\in\mathcal{E}_{w,\epsilon}(\psi)$ but one still has $p\in E\cdot w^{-1}q$. So if we construct an open cover of $\mathcal{E}_{w,\epsilon}(\psi)$ directly from cutting all the subsets of the form $E\cdot w^{-1}q$ into cubes, there would be many unnecessary open cubes counted in the open cover. In order to obtain an upper bound of the Hausdorff dimension of $\mathcal{E}_{w,\epsilon}(\psi)$, we project the subset $\mathcal{E}_{w,\epsilon}(\psi)$ into the compact quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap\Gamma)$, and construct an open cover for the projection of $\mathcal{E}_{w,\epsilon}(\psi)$ instead. The projections of the subsets of the form $E\cdot w^{-1}q$ overlap in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap\Gamma)$, and by a careful analysis, one can show that the union of these projections can be described in a quite economical way, from which an efficient open cover can be constructed for the projection of $\mathcal{E}_{w,\epsilon}(\psi)$. Then by the countable stability of Hausdorff dimension, we may obtain an upper bound of $\dim_{H}\mathcal{E}_{w,\epsilon}(\psi)$ for each $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, and consequently an upper bound for $\dim_{H}S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Note that in this way, we are able to avoid many computations in [25], and the only important information we need is about the growth rates of the volume estimates of certain real algebraic varieties and the growth rates of the numbers of rational points in these algebraic varieties, for which we may appeal to algebraic geometry (Cf. Manin’s conjecture [3]).
The paper is organized as follows:
  1. $\bullet$
    In §2, we first give introductions to the reduction theory of arithmetic lattices and the theory of irreducible representations of complex semisimple groups and complex Lie algebras. Then we prove Theorem 1.3. After that, we define a notion of rational elements and discuss some properties of the rational elements which will be important in the subsequent sections.
  2. $\bullet$
    In §3, we assume that the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on any arithmetic quotient $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is mixing, where $\mathbf{G}(\mathbb{R})^{0}$ is the identity component of $\mathbf{G}(\mathbb{R})$ (as a Lie group) and $\Gamma$ is a subgroup of finite index in $\mathbf{G}(\mathbb{Z})$. We select certain rational elements in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ according to some algebraic conditions, and apply the mixing property to obtain an asymptotic estimate about the measure of the subset of the selected rational elements. In §4, we construct a Cantor-type subset $\mathbf{A}_{\infty}$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ from neighborhoods of these selected rational elements. With the asymptotic estimate result established in §3 and the condition that $0\leq\tau(\psi)<\beta_{0}(a_{-1})$, we compute the Hausdorff dimension of $\mathbf{A}_{\infty}$ and prove Theorem 1.4.
  3. $\bullet$
    In §5, we discuss the case where $0\leq\tau(\psi)<\beta_{0}(a_{-1})$ but the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on any arithmetic quotient of $\mathbf{G}(\mathbb{R})^{0}$ is not mixing. Then there is a proper normal $\mathbb{Q}$-subgroup $\mathbf{H}$ of $\mathbf{G}$ such that $\mathbf{H}$ contains $\{a_{t}\}_{t\in\mathbb{R}}$ and the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on any arithmetic quotient of $\mathbf{H}(\mathbb{R})^{0}$(= the identity component of $\mathbf{H}(\mathbb{R})$) is mixing. We show that the arguments in §3 and §4 work equally well in the group $\mathbf{H}(\mathbb{R})$ and Theorem 1.4 still holds in this case. At the end of this section, we discuss the case $\tau(\psi)=\beta_{0}(a_{-1})$. This completes the proof of Theorem 1.4.
  4. $\bullet$
    In §6, we divide the subset $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ into finitely many subsets $\mathcal{E}_{w,\epsilon}(\psi)$ indexed by elements $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$, each of which corresponds to a type of Diophantine approximation. We project the subsets $\mathcal{E}_{w,\epsilon}(\psi)$ into the quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap\Gamma$ and construct open covers for the projections of $\mathcal{E}_{w,\epsilon}(\psi)$. This gives an upper bound of the Hausdorff dimension of $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ and proves Theorem 1.5.
  5. $\bullet$
    In §7, we derive Theorems 1.61.7 and 1.8 from Theorems 1.4 and  1.5. The proofs of Theorem 1.12 and Theorem 1.15 are given in §8.
@@ 2. Preliminaries
In this section, we list some preliminaries needed in this paper and prove Theorem 1.3. Then we give a definition of rational elements in $\mathbf{G}(\mathbb{R})$ and discuss some properties of rational elements. Let $\mathbf{G}$, $\mathbf{T}$, $\{a_{t}\}_{t\in\mathbb{R}}$ and the representation $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ satisfy the assumption in §1.2.
We first need the reduction theory of arithmetic subgroups of $\mathbf{G}(\mathbb{R})$ [9, 7]. Let $K$ be a maximal compact subgroup in $\mathbf{G}(\mathbb{R})$ and $\Gamma$ an arithmetic subgroup in $\mathbf{G}(\mathbb{Z})$. We fix a minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ in $\mathbf{G}$ containing $\mathbf{T}$ with the Levi subgroup $Z(\mathbf{T})$ (the centralizer of $\mathbf{T}$ in $\mathbf{G}$). Denote by $\Phi$ the set of $\mathbb{Q}$-roots in $\mathbf{G}$ with respect to $\mathbf{T}$, $\Phi^{+}$ the set of positive $\mathbb{Q}$-roots corresponding to the minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ and $\Delta$ the set of simple $\mathbb{Q}$-roots in $\Phi^{+}$. For any parabolic $\mathbb{Q}$-subgroup $\mathbf{P}$, denote by $R_{u}(\mathbf{P})$ its unipotent radical. We write $\mathbf{P}_{0}=\mathbf{M}_{0}\cdot R_{u}(\mathbf{P}_{0})$ where $\mathbf{M}_{0}=Z(\mathbf{T})$ is the Levi subgroup, and write $\mathbf{M}_{a}$ for the maximal $\mathbb{Q}$-anisotropic subgroup in $\mathbf{M}_{0}$ so that $\mathbf{M}_{0}=\mathbf{T}\cdot\mathbf{M}_{a}$. Let $M=\mathbf{M}_{a}(\mathbb{R})^{0}$ be the identity component of the Lie group $\mathbf{M}_{a}(\mathbb{R})$. For $\eta>0$, denote by
@@
A Siegel set in $\mathbf{G}(\mathbb{R})$ is a subset of the form $S_{\eta,\Omega}=K\cdot T_{\eta}\cdot\Omega$ for some $\eta>0$ and some relatively compact open subset $\Omega$ containing identity in $M\cdot R_{u}(\mathbf{P}_{0})(\mathbb{R})$, and the group $\mathbf{G}(\mathbb{R})$ can be written as
@@
for some Siegel set $S_{\eta,\Omega}$ and a finite subset $\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$. Moreover, the finite set $\mathcal{K}$ satisfies the property that
@@
where $\mathbf{P}_{0}$ is the minimal parabolic $\mathbb{Q}$-subgroup in $\mathbf{G}$. Denote by
@@
and we may assume that $e\in\mathcal{K}$. In what follows, we choose $\Gamma$ to be an arithmetic subgroup in $\mathbf{G}(\mathbb{Z})\cap\mathbf{G}(\mathbb{R})^{0}$, where $\mathbf{G}(\mathbb{R})^{0}$ denotes the identity component of the Lie group $\mathbf{G}(\mathbb{R})$, so that $\rho(\Gamma)$ preserves the lattice $\mathbb{Z}^{d}$ in $V$. Without loss of generality, we may assume that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\mathbf{P}_{0})$ and the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\overline{\mathbf{P}}_{0})$, where $\overline{\mathbf{P}}_{0}$ is the opposite minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{0}$ determined by $\Phi\setminus\Phi^{+}$ with the same Levi subgroup $\mathbf{M}_{0}=Z(\mathbf{T})$.
Now we choose a maximal $\mathbb{Q}$-torus $\mathbf{S}$ in $\mathbf{P}_{0}$ containing $\mathbf{T}$. The Lie algebra $\mathfrak{g}$ of $\mathbf{G}$ can be written as a direct sum of root spaces relative to $\mathbf{S}$ via the adjoint representation of $\mathbf{G}$
$\mathfrak{g}=\mathfrak{g}_{0}\oplus\bigoplus_{\alpha\in\Psi}\mathfrak{g}_{\alpha},$
where $\Psi$ is the set of roots in $\mathfrak{g}$ relative to $\mathbf{S}$. We may determine a set of positive roots in $\Psi$, which we denote by $\Psi^{+}$, such that
@@
Here $\operatorname{Lie}(R_{u}(\mathbf{P}_{0}))$ and $\operatorname{Lie}(R_{u}(\overline{\mathbf{P}}_{0}))$ denote the Lie algebras of $R_{u}(\mathbf{P}_{0})$ and $R_{u}(\overline{\mathbf{P}}_{0})$ respectively. The set of simple roots in $\Psi^{+}$ is denoted by $\Pi$.
It is known that there are complete classifications of irreducible representations of complex semisimple groups and Lie algebras [41, 55]. Let $V$ be the vector space in the representation $\rho$, and
$V=\bigoplus_{\lambda}V_{\lambda}$
the decomposition of $V$ into the direct sum of weight spaces $V_{\lambda}$ relative to $\mathbf{S}$. According to the theorem of highest weight, there is a unique highest weight $\lambda_{0}$ among the weights $\lambda$’s (the order is determined by $(\Psi,\Psi^{+},\Pi)$) such that